What is the product AB of the following complex numbers,in polar form?
A=3+4i B=8+6i

Answer by Alan3354(69443)

What is the product AB of the following complex numbers,in polar form?
A=3+4i B=8+6i
----------------
A*B = 24+18i + 32i-24 = 50i
= 50cis(90)